OpenStudy (anonymous):

MEDALS AWARDED eliminate the parameter. x = square root of t y = 2t + 5

OpenStudy (aravindg):

\[x=\sqrt{ty}\] Square it and solve for t

OpenStudy (anonymous):

Okay but I need a y = equation @AravindG

OpenStudy (aravindg):

Once you can get t substitute in the equation.

OpenStudy (anonymous):

Options: a. y = 2 sqrt x+ 5, x ≥ 0 b. y = 2 sqrt x - 5, x ≥ 0 c. y = 2x2 + 5 d. y = 2x2 - 5

OpenStudy (anonymous):

How do I find t?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

x = sqrt t square both sides u will get t now plug this t in your y= ... equation

OpenStudy (aravindg):

@hartnn x = square root of t y

hartnn (hartnn):

not actually its the way she has written, its confusing actually x= sqrt t and y=2t+5

OpenStudy (aravindg):

oh ..Please use eqn editor next time

hartnn (hartnn):

looking at tue choices, I m pretty sure about that

OpenStudy (anonymous):

Sorry. But yeah, @hartnn is right about how it's written.

hartnn (hartnn):

did u get t in terms of x?

OpenStudy (anonymous):

Ha okay, wait. The answer is c (y = 2x2 + 5 ). Right?

hartnn (hartnn):

yes:)

OpenStudy (anonymous):

Thank you :)